### v3

# DriveDreamer4D: World Models Are Effective Data Machines for 4D Driving Scene Representation

Abstract
Closed-loop simulation is essential for advancing end-to-end autonomous driving systems. Contemporary sensor simulation methods, such as NeRF and 3DGS, rely predominantly on conditions closely aligned with training data distributions, which are largely confined to forward-driving scenarios. Consequently, these methods face limitations when rendering complex maneuvers (e.g., lane change, acceleration, deceleration). Recent advancements in autonomous-driving world models have demonstrated the potential to generate diverse driving videos. However, these approaches remain constrained to 2D video generation, inherently lacking the spatiotemporal coherence required to capture intricacies of dynamic driving environments. In this paper, we introduce DriveDreamer4D, which enhances 4D driving scene representation leveraging world model priors. Specifically, we utilize the world model as a data machine to synthesize novel trajectory videos, where structured conditions are explicitly leveraged to control the spatial-temporal consistency of traffic elements. Besides, the cousin data training strategy is proposed to facilitate merging real and synthetic data for optimizing 4DGS. To our knowledge, DriveDreamer4D is the first to utilize video generation models for improving 4D reconstruction in driving scenarios. Experimental results reveal that DriveDreamer4D significantly enhances generation quality under novel trajectory views, achieving a relative improvement in FID by 32.1%, 46.4%, and 16.3% compared to PVG, $\text{S}^{3}$aussian, and Deformable-GS. Moreover, DriveDreamer4D markedly enhances the spatiotemporal coherence of driving agents, which is verified by a comprehensive user study and the relative increases of 22.6%, 43.5%, and 15.6% in the NTA-IoU metric.
Figure 1: Previous 4D Gaussian Splatting methods (e.g., PVG [9], $\text{S}^{3}\text{Gaussian}$ [28], Deformable-GS [74]) face challenges in rendering novel trajectories, such as lane change. DriveDreamer4D addresses this by enhancing 4D driving scene representation via integrating priors from world models, significantly improving rendering quality under complex scenarios and novel trajectory viewpoints.
1 Introduction
End-to-end planning [26, 27, 30], which directly maps sensor inputs to control signals, is among the most critical and promising tasks in autonomous driving. However, current open-loop evaluations are inadequate for accurately assessing end-to-end planning algorithms, highlighting an urgent need for enhanced evaluation methods [39, 80, 37]. A compelling solution lies in closed-loop evaluations within real-world scenarios, which require retrieving sensor data from arbitrarily specified viewpoints. This necessitates constructing a 4D driving scene representation capable of reconstructing complex, dynamic driving environments.
Closed-loop simulation in driving environments predominantly relies on scene reconstruction techniques such as Neural Radiance Fields (NeRF) [45, 71, 73, 18] and 3D Gaussian Splatting (3DGS) [32, 70, 28, 11], which are inherently limited by the density of input data. Specifically, these methods can render scenes effectively only under conditions closely aligned with their training data distributions—primarily forward-driving scenarios—and struggle to perform accurately during complex maneuvers (see Fig. 1). To mitigate these limitations, methods like SGD [78] and GGS [20] leverage generative models to extend the range of training viewpoints. However, these approaches primarily supplement sparse image data or static background elements, falling short of modeling the intricacies of dynamic, interactive driving scenes. Recently, advancements in autonomous driving world models [61, 81, 63, 25, 64, 16] have introduced the capability to generate diverse, command-aligned video viewpoints, offering renewed promise for closed-loop simulation in autonomous driving. Nonetheless, these models remain constrained to 2D videos, lacking the spatial-temporal coherence essential for accurately modeling complex driving scenarios.
In this paper, we introduce DriveDreamer4D, which improves 4D driving scene representation by integrating priors from autonomous driving world models. Our approach utilizes an autonomous driving world model [81] as a generative engine, synthesizing novel trajectory video data that densifies real-world driving datasets for enhanced training. Notably, we propose the Novel Trajectory Generation Module (NTGM) to generate diverse structured traffic conditions, and DriveDreamer4D applies these conditions to independently regulate the motion dynamics of foreground and background elements in complex driving environments. These conditions undergo view projection synchronized with vehicle maneuvers, ensuring that the synthesized data adheres to the spatiotemporal constraints. Subsequently, the Cousin Data Training Strategy (CDTS) is proposed to merge temporal-aligned real and synthetic data for training 4DGS. In CDTS, a regularization loss is further incorporated to ensure perceptual coherence. To the best of our knowledge, DriveDreamer4D is the first framework to harness video generation models for elevating 4D scene reconstruction quality in autonomous driving, providing richly varied viewpoint data for scenarios including lane change, acceleration, and deceleration. As shown in Fig. 1, experiment results demonstrate that DriveDreamer4D significantly enhances generation fidelity for novel trajectory viewpoints, achieving a relative improvement in FID by 32.1%, 46.4%, and 16.3% compared to PVG [9], $\text{S}^{3}$aussian [28], and Deformable-GS [74]. Besides, DriveDreamer4D fortifies the spatiotemporal coherence between foreground and background elements, with respective increases of 22.6%, 43.5%, and 15.6% in the NTA-IoU metric. Furthermore, a comprehensive user study confirms that the average win rate of DriveDreamer4D exceeds 80%, compared to three baselines.
The primary contributions of this work are as follows: (1) We present DriveDreamer4D, the first framework to leverage world model priors for advancing 4D scene reconstruction in autonomous driving. (2) The NTGM is proposed to automate the generation of structured conditions, allowing DriveDreamer4D to create novel trajectory videos with complex maneuvers while ensuring spatial-temporal consistency. Additionally, the CDTS is introduced to merge temporal-aligned real and synthetic data for training 4DGS, using a regularization loss to maintain perceptual coherence. (3) We perform comprehensive experiments to validate that DriveDreamer4D notably enhances generation quality across novel trajectory viewpoints, as well as the spatiotemporal coherence of driving scene elements.
2 Related Work
2.1 Driving Scene Representation
NeRF and 3DGS have emerged as leading approaches for 3D scene representation. NeRF models [45, 2, 3, 46] continuous volumetric scenes using multi-layer perceptron (MLP) networks, enabling highly detailed scene reconstructions with remarkable rendering quality. More recently, 3DGS [32, 77] introduces an innovative method by defining a set of anisotropic Gaussians in 3D space, leveraging adaptive density control to achieve high-quality renderings from sparse point cloud inputs. Several works have extended NeRF [73, 71, 29, 43, 52, 58, 18, 12] or 3DGS [70, 82, 9, 78, 28, 11, 40] to autonomous driving scenarios. Given the dynamic nature of driving environments, there has also been significant effort in modeling 4D driving scene representations. Some approaches encode time as an additional input to parameterize 4D scenes [1, 13, 38, 42, 48, 56, 28], while others represent scenes as a composition of moving object models alongside a static background model [73, 35, 47, 59, 66, 68]. Despite these advancements, methods based on NeRF and 3DGS face limitations tied to the density of input data. These techniques can only render scenes effectively when sensor data closely matches the training data distribution, which is typically confined to forward-driving scenarios.
2.2 World Models
The world model module predicts possible future world states as a function of imagined action sequences proposed by the actor [36, 83]. Approaches such as [63, 69, 79, 17, 19, 4, 22, 44, 23, 5, 34, 75, 24, 67, 62] simulate environments through video generation controlled by free-text actions. At the forefront of this evolution is Sora [6], which leverages advanced generative techniques to produce intricate visual sequences that respect the fundamental laws of physics. This ability to deeply understand and simulate the environment not only improves video generation quality but also has substantial implications for real-world driving scenarios. Autonomous driving world models [61, 81, 64, 25, 16, 72] employ predictive methodologies to interpret driving environments, thereby generating realistic driving scenarios and learning key driving elements and policies from video data. Although these models successfully produce diverse driving video data conditioned on complex driving actions, they remain limited to 2D outputs and lack the spatial-temporal coherence needed to accurately capture the complexities of dynamic driving environments.
2.3 Diffusion Prior for 3D Representation
Constructing comprehensive 3D scenes from limited observations demands generative prior, particularly for unseen areas. Earlier studies distill the knowledge from text-to-image diffusion models [53, 49, 54, 51] into a 3D representation model. Specifically, the Score Distillation Sampling (SDS) [50, 41, 65] is adopted to synthesize a 3D object from the text prompt. Furthermore, to enhance 3D consistency, several approaches extend the multi-view diffusion models [55, 15] and video diffusion models [4, 60, 10] to 3D scene generation. To extend the diffusion prior to complex, dynamic, large-scale driving scenes for 3D reconstruction, methods such as SGD [78], GGS [20] and MagicDrive3D [14] employ generative models to broaden the range of training viewpoints. Nonetheless, these approaches mainly address sparse image data or static background elements, lacking the capacity to fully capture the complexities inherent in the 4D driving environments.
Figure 2: The overall framework of DriveDreamer4D. Initially, by altering the actions of the original trajectory (e.g., steering angle, speed), new trajectories can be obtained. Conditioned on the first frame and the structured information (3D bounding boxes, HDMap) from the new trajectory, the novel trajectory videos are generated. Subsequently, the temporal-aligned cousin pair (original and novel trajectory videos) are merged to optimize the 4D Gaussian Splatting model, where a regularization loss is calculated to ensure perceptual coherence.
3 Method
In this section, we first elaborate on the preliminaries of 4D driving scene representation and world models for driving video generation. Then we present the details of DriveDreamer4D, which enhances 4D driving scene representation leveraging priors from driving world models.
3.1 Preliminary
3.1.1 4D Driving Scene Representation
4DGS models the driving scene with a collection of 3DGS and a temporal field module. Each 3DGS [32] is parameterized by its center position $\boldsymbol{x}$, opacity $\boldsymbol{\gamma}$, covariance $\boldsymbol{\Sigma}$, and view-dependent RGB color $\boldsymbol{c}$, controlled via spherical harmonics. For stability, each covariance matrix $\boldsymbol{\Sigma}$ is decomposed by:
${\boldsymbol{\Sigma}}={\boldsymbol{RSS}}^{T}{\boldsymbol{R}}^{T},$ (1)
where scaling matrix $\boldsymbol{S}$ and a rotation matrix $\boldsymbol{R}$ are learnable parameters, represented by scaling $\boldsymbol{s}$ and quaternion $\boldsymbol{r}$. All trainable parameters of a single 3D Gaussian are collectively denoted as $\boldsymbol{\phi}=\{\boldsymbol{x,\gamma,s,r,c}\}$. The temporal field $\mathcal{F}$ takes $\boldsymbol{\phi}$ and a time step $\boldsymbol{t}_{gs}$ as input, outputting the offset $\delta\boldsymbol{\phi}=\{\delta\boldsymbol{x},\delta\boldsymbol{\gamma},\delta\boldsymbol{s},\delta\boldsymbol{r},\delta\boldsymbol{c}\}$ for each Gaussian relative to canonical space. The 4D Gaussian $\boldsymbol{\phi}^{\prime}=\{\boldsymbol{x}^{\prime},\boldsymbol{\gamma}^{\prime},\boldsymbol{s}^{\prime},\boldsymbol{r}^{\prime},\boldsymbol{c}^{\prime}\}$ is then computed by:
${\boldsymbol{\phi}}^{\prime}={\boldsymbol{\phi}}+\delta{\boldsymbol{\phi}}={\boldsymbol{\phi}}+\mathcal{F}({\boldsymbol{\phi},{\boldsymbol{t}_{gs}}}).$ (2)
Following [76], a differentiable Gaussian Splatting renderer is employed to project 4D Gaussians $\boldsymbol{\phi}$ into camera coordinates, yielding the covariance matrix ${\boldsymbol{\Sigma}^{\prime}}={\boldsymbol{JV\Sigma V}}^{T}{\boldsymbol{J}}^{T}$, where ${\boldsymbol{J}}$ is the Jacobian matrix of the perspective projection, and ${\boldsymbol{V}}$ is the transform matrix. The color of each pixel is calculated by $N$ ordered points using $\alpha$-blending:
$C=\sum_{i\in N}T_{i}{\boldsymbol{c}^{\prime}_{i}}{\alpha_{i}},$ (3)
where $T_{i}$ is the transmittance defined by $\prod_{j=1}^{i-1}(1-{\alpha_{j}})$, $\boldsymbol{c}_{i}^{\prime}$ denotes the color of each point, $\alpha_{i}$ is given by evaluating a 2D Gaussian with covariance $\boldsymbol{\Sigma}^{\prime}$ multiplied with a learned per-point opacity $\boldsymbol{\gamma}_{i}^{\prime}$. The trainable parameters $\boldsymbol{\phi}^{\prime}$ can be optimized by a combination of RGB loss, depth loss and SSIM loss:
$\mathcal{L}_{\text{ori}}(\phi^{\prime})$ $=\lambda_{1}\|\hat{I}_{\text{ori}}-I_{\text{ori}}\|_{1}+\lambda_{2}\|\hat{D}_{\text{ori}}-D_{\text{ori}}\|_{1}$ (4)
$+\lambda_{3}\textrm{SSIM}(\hat{I}_{\text{ori}},I_{\text{ori}}),$
where $\hat{I}_{\text{ori}}$ and $I_{\text{ori}}$ represent the rendered image and the ground truth image. $\hat{D}_{\text{ori}}$ and $D_{\text{ori}}$ are the rendered depth and the ground truth LiDAR depth map. $\textrm{SSIM}(\cdot)$ refers to the operation of the Structural Similarity Index Measure, and $\lambda_{1},\lambda_{2},\lambda_{3}$ are the loss weights.
3.1.2 World Models for Driving Video Generation
The world model module predicts possible future world states based on imagined action sequences [36]. Autonomous-driving world models [61, 81, 64, 16], typically based on diffusion models, leverage structured driving information or action controls to guide future video prediction. During training, these models first encode videos $\boldsymbol{v}$ into a lower-dimensional latent space $\boldsymbol{z}=\mathcal{E}(\boldsymbol{v})$ using a variational encoder $\mathcal{E}$. After adding noise $\epsilon_{t}$ to the latent, the diffusion model learns a denoising process. This diffusion process is optimized by:
$\mathcal{L}_{diff}=\mathbb{E}_{\boldsymbol{z},\epsilon\sim\mathcal{N}(0,1),t}\left[\left\|\epsilon_{t}-\epsilon_{\theta}\left(\boldsymbol{z}_{t},t,\boldsymbol{f}\right)\right\|_{2}^{2}\right],$ (5)
where $\epsilon_{\theta}$ is a parameterized denoising network, $t$ denotes the time step, representing the level of noise added or removed at each stage. Additionally, to improve the controllability of the generated data, conditional features $\boldsymbol{f}$ (e.g., reference images, speed, steering angle, scene layouts, camera poses and textual information) can be introduced into the reverse diffusion process, ensuring that the generated outputs adhere to the input control signals. During inference, the world models can be conditioned on a reference image to control the style of the output scene, while predicting the future world states contingent upon the other input actions.
3.2 DriveDreamer4D
The overall pipeline of DriveDreamer4D is depicted in Fig. 2. In the upper part, the Novel Trajectory Generation Module (NTGM) is proposed to adjust driving actions (e.g., steering angle, speed) to generate new trajectories. These novel trajectories provide new perspectives for extracting structured information like 3D boxes and HDMap. Subsequently, a controllable video diffusion model synthesizes videos from these updated viewpoints, incorporating specific priors associated with the modified trajectories. In the lower part, the Cousin Data Training Strategy (CDTS) is introduced to combine the temporal-aligned original and generated data for optimizing the 4DGS model, where a regularization loss is calculated to impose perceptual coherence. In the following sections, we delve into the details of novel trajectories video generation and then introduce the CDTS for 4D reconstruction.
3.2.1 Novel Trajectory Video Generation
As previously mentioned, traditional 4DGS methods are limited in rendering complex maneuvers, largely due to the training data being dominated by straightforward driving scenarios. To overcome this, DriveDreamer4D leverages world model priors to generate diverse viewpoint data, enhancing the 4D scene representation. To achieve this, we propose the NTGM, which is designed to create new trajectories that serve as input for the world model, enabling the automated generation of complex maneuver data. NTGM comprises two main components: (1) novel trajectory proposal, (2) trajectory safety assessment. In the novel trajectory proposal stage, text-to-trajectory [81] can be adopted to automatically generate diverse complex trajectories. Additionally, trajectories can be custom-designed to meet specific requirements, allowing for tailored data generation based on precise needs. The overview of the custom-designed trajectory proposal (e.g., lane change) and trajectory safety assessment is shown in the Algo. 1. In a specific driving scenario, the original trajectory in the world coordinate system can be readily acquired as $\mathcal{T}_{\text{ori}}^{\text{world}}=\{p_{i}^{\text{world}}\}_{i=0}^{K}$, where $K$ denotes the number of frames and $p_{i}^{\text{world}}\in\mathbb{R}^{3}$ refers to the position of the ego-vehicle at the $i$-th frame. To propose novel trajectories, the original trajectory $\mathcal{T}_{\text{ori}}^{\text{world}}$ is transformed into the ego-vehicle coordinate system of the first frame, denoted as $\mathcal{T}_{\text{ori}}^{\text{EgoStart}}$ and computed as:
=${}^{T}=M_{0}^{-1}\times[p_{i}^{\text{world}},1]^{T},$ (6)
where $M_{0}\in\mathbb{R}^{4\times 4}$ represents the transformation matrix from the ego-vehicle coordinate system of the first frame to the world coordinate system, []delimited-[][$\cdot$][ ⋅ ] denotes the operation of the concat. In the ego-vehicle coordinate system, the vehicle’s heading is aligned with the positive $x$-axis, the $y$-axis points to the left side of the vehicle, and the $z$-axis is oriented vertically upwards, perpendicular to the plane of the vehicle. Consequently, changes in the vehicle’s velocity and direction can be respectively represented by adjusting the value along the $x$-axis and $y$-axis. A final safety assessment is conducted for the newly generated trajectory points, which includes verifying whether the vehicle trajectories $p$ remain within drivable areas $\mathcal{B}_{\text{road}}$ and ensuring that no collisions occur with pedestrians or other vehicles $\{o_{j}\}_{j=1}^{M}$.
$\begin{array}[]{ll}p\in\mathcal{B}_{\text{road}},&\\[8.0pt] \|p-o_{j}\|\geq d_{\text{min}},&\quad\forall j\in\{1,\dots,M\},\end{array}$ (7)
where $d_{\text{min}}$ is the minimal distance between different agents. Once a novel trajectory that complies with traffic regulations is generated, the road structure and 3D bounding boxes can be projected onto the camera view from the perspective of the new trajectory, thereby generating structured information relative to the updated trajectory. This structured information, along with the initial frame and text, is fed into a world model [81] to produce the videos that follow the novel trajectories.
Algorithm 1 Novel Trajectory Generation Module
0:  Trajectory $\mathcal{T}_{\text{ori}}^{\text{world}}$, Transformation matrix $M_{0}$
0:  Novel trajectory $\mathcal{T}_{\text{novel}}^{\text{ego}}$
  $\mathcal{T}_{\text{novel}}^{\text{ego}}\leftarrow[[0,0,0]]$
  Offset $\leftarrow$ 0
  for each $p^{\text{world}}_{\text{{ori}}}$ in $\mathcal{T}_{\text{ori}}^{\text{world}}$[1:] do
     $p^{\text{Ego}\text{Start}}\leftarrow$ RelativeCoord($p^{\text{world}}_{\text{{ori}}}$, $M_{0}$)
     MaxOffset $\leftarrow$ 0.1
     while True do
        NewOffset $\leftarrow$ Offset + RandOffset(0, MaxOffset)
        $p^{\text{Ego}\text{Start}^{\prime}}\leftarrow p^{\text{Ego}\text{Start}}$ + [0, NewOffset, 0]
        if SafeCheck($p^{\text{EgoStart}^{\prime}}$then
           AddElement($\mathcal{T}_{\text{novel}}^{\text{ego}}$, $p^{\text{EgoStart}^{\prime}}$)
           Offset $\leftarrow$ NewOffset
           break
        else
           MaxOffset $\leftarrow$ MaxOffset/2
        end if
     end while
  end for
Method Lane Change Acceleration Deceleration Average
NTA-IoU $\uparrow$ NTL-IoU $\uparrow$ NTA-IoU $\uparrow$ NTL-IoU $\uparrow$ NTA-IoU $\uparrow$ NTL-IoU $\uparrow$ NTA-IoU $\uparrow$ NTL-IoU $\uparrow$
PVG [9] 0.256 50.70 0.396 53.08 0.394 53.65 0.349 52.48
DriveDreamer4D with PVG 0.438 53.06 0.421 53.35 0.424 53.89 0.428 53.43
$\text{S}^{3}$aussian [28] 0.175 49.05 0.434 51.93 0.384 52.14 0.331 51.04
DriveDreamer4D with $\text{S}^{3}$aussian 0.495 53.42 0.484 52.63 0.445 52.69 0.475 52.91
Deformable-GS [74] 0.240 51.62 0.346 52.17 0.377 53.21 0.321 52.33
DriveDreamer4D with Deformable-GS 0.335 52.93 0.371 52.77 0.406 53.79 0.371 53.16
Table 1: Comparison of NTA-IoU and NTL-IoU scores across different novel trajectory views (lane change, acceleration, deceleration).
3.2.2 Cousin Data Training Strategy
To better integrate generated data for training 4DGS, we propose the CDTS. Specifically, we construct temporally aligned cousin pair data as a minimal training batch:
$\text{BatchStack}(\{\hat{I}_{\text{ori},t}\}_{t=0}^{T},\{\hat{I}_{\text{novel},t}\}_{t=0}^{T}),$ (8)
where $\text{BatchStack}(\cdot)$ is the data processor to stack temporal-aligned $\{\hat{I}_{\text{ori},t}\}_{t=0}^{T}$ and $\{\hat{I}_{\text{novel},t}\}_{t=0}^{T}$ into a training batch. By leveraging real and synthetic data aligned at each timestep, CDTS mitigates the data gap in 4DGS training, enhancing the model’s ability to learn consistent representations across real and synthetic data. To optimize 4DGS, the temporal-aligned cousin pair is input per step before gradient optimization. The loss function $\mathcal{L}_{\text{ori}}$ for the original data is defined in Eq. 4. And the loss function $\mathcal{L}_{\text{novel}}$ for the generated data is akin to [9, 28], defined as follows:
$\mathcal{L}_{\text{novel}}(\phi^{\prime})$ $=\lambda_{\text{1}}\|\hat{I}_{\text{novel}}-I_{\text{novel}}\|_{1}$ (9)
$\quad+\lambda_{\text{3}}\textrm{SSIM}(\hat{I}_{\text{novel}},I_{\text{novel}}),$
where $I_{\text{novel}}$ represents the generated images corresponding to the novel trajectories as described in Sec. 3.2.1, and $\hat{I}_{\text{novel}}$ denotes the rendered images under the novel trajectories via differentiable splatting [76]. Notably, different from [9, 28], depth maps are not employed as constraints in the optimization of 4DGS when using the generated dataset $D_{\text{novel}}$. The limitation arises from the fact that LiDAR point cloud data is exclusively collected for the original trajectory. When these LiDAR points are projected onto a new trajectory, it cannot produce a complete depth map for the new perspective, as something visible in the novel trajectory may have been occluded in the original view. Consequently, the incorporation of such depth maps does not facilitate the optimization of the 4DGS model. More details are described in Sec. 4.3. Additionally, we propose a regularization loss to enhance the perceptual coherence, defined as follows:
$\mathcal{L}_{\text{reg}}(\phi^{\prime})=\|\mathcal{F}_{p}(\hat{I}_{\text{ori}})-\mathcal{F}_{p}(\hat{I}_{\text{novel}})\|_{1},$ (10)
where $\mathcal{F}_{p}(\cdot)$ denotes the perception feature extraction model [21]. The overall loss function for mixed training is defined as follows:
$\mathcal{L}(\phi^{\prime})=\mathcal{L}_{\text{ori}}+\lambda_{\text{novel}}\mathcal{L}_{\text{novel}}+\lambda_{\text{reg}}\mathcal{L}_{\text{reg}}.$ (11)
4 Experiments
In this section, we first outline the experimental setup, including the dataset, implementation details, and evaluation metrics. Subsequently, both quantitative and qualitative evidence are provided to demonstrate that the proposed DriveDreamer4D significantly enhances rendering quality for novel trajectory viewpoints and improves the spatiotemporal coherence of foreground and background components. Finally, we conduct an ablation study on the hyperparameter settings, as well as the effects of depth loss and the proposed CDTS, including temporal-aligned pairs, and regularization loss.
4.1 Experiment Setup
Dataset. We conduct experiments using the Waymo dataset [57], known for its comprehensive real-world driving logs. However, most logs capture scenes with relatively straightforward dynamics, lacking focus on scenarios with dense, complex vehicle interactions. To address this gap, we specifically select eight scenes characterized by highly dynamic interactions, featuring numerous vehicles with diverse relative positions and intricate driving trajectories. Each selected segment contains approximately 40 frames, with segment IDs detailed in the supplement.
Implementation Details. To demonstrate the versatility and robustness of DriveDreamer4D, we incorporate various 4DGS baselines into our pipeline, including Deformable-GS [74], $\text{S}^{3}$aussian [28], and PVG [9]. For a fair comparison, LiDAR supervision is introduced to Deformable-GS. During training, scenes are segmented into multiple clips, each containing 40 frames, aligned with the generative model’s output length. We use only forward-facing camera data and standardize the resolution across methods to $640\times 960$. Our models are trained using the Adam optimizer [33], following the learning rate schedule used for 3D Gaussian Splatting [32]. Hyperparameter settings are aligned with each baseline [74, 28, 9], and the training strategy remains the same, with the exception of the incorporation of CDTS.
Figure 3: Qualitative comparisons of novel trajectory renderings during lane change scenarios. The orange boxes highlight that DriveDreamer4D significantly enhances the rendering quality across various baselines (PVG [9], $\text{S}^{3}$aussian [28], Deformable-GS [74]).
Metrics. Traditional 3D reconstruction tasks typically employ PSNR and SSIM metrics for evaluation, with validation sets that closely match the training data distribution (i.e., uniformly sampling frames from video sequences for validation, with the remainder used for training). However, in closed-loop driving simulation, the focus shifts to evaluating model rendering performance under novel trajectories, where corresponding sensor data are unavailable, making metrics like PSNR and SSIM inapplicable for evaluation. Therefore, we propose Novel Trajectory Agent IoU (NTA-IoU) and Novel Trajectory Lane IoU (NTL-IoU), which assess the spatiotemporal coherence of foreground and background traffic components in novel trajectory viewpoints.
Method FID $\downarrow$
PVG[9] 105.29
DriveDreamer4D with PVG 71.52
$\text{S}^{3}$aussian[28] 124.90
DriveDreamer4D with $\text{S}^{3}$aussian 66.93
Deformable-GS[74] 92.34
DriveDreamer4D with Deformable-GS 77.32
Table 2: Comparison of FID scores in novel trajectory view synthesis (lane change) on the Waymo dataset.
For NTA-IoU, we use YOLO11 [31] to identify vehicles in images rendered from novel trajectory views, yielding 2D bounding boxes. Simultaneously, geometric transformations are applied to the original 3D bounding boxes, projecting them onto the new viewpoints to generate corresponding 2D bounding boxes. For each projected 2D box, we then identify the closest detector-generated 2D box and compute their Intersection over Union (IoU). To ensure accurate matching, a distance threshold $d_{\text{thresh}}$ is introduced: when the center-to-center distance $\|c(B^{\text{proj}})-c(B^{\text{det}})\|$ between the nearest detected box $B^{\text{det}}$ and the correctly projected box $B^{\text{proj}}$ surpasses this threshold, their NTA-IoU is assigned a value of zero:
[…]
For NTL-IoU, we employ TwinLiteNet [8] to extract 2D lanes from rendered images. Ground truth lanes are also projected onto the 2D image plane. We then compute the mean Intersection-over-Union (mIoU) between the rendered and ground truth lanes $L^{\text{det}}$ and $L^{\text{proj}}$:
[…]
Additionally, in lane change scenarios, we observe inaccuracies in relative positioning, as well as frequent occurrences of artifacts such as flying points and ghosting, which notably degrade image quality. To assess this, we employ the FID metric [21], which quantifies differences in feature distribution between rendered novel trajectory images and original trajectory images. This metric effectively reflects visual quality and is particularly sensitive to artifacts like flying points and ghosting, providing a robust measure of image fidelity in these complex scenes. Finally, a user study is conducted to evaluate the quality of the renderings, where participants compare the rendering results of each baseline with its DriveDreamer4D enhanced version across three novel trajectories. The evaluation criteria focus on overall video quality, with particular attention to foreground objects like vehicles. For each comparison, participants were asked to select the option they found most favorable. Further details are provided in the supplement.
Counterpart Method DriveDreamer4D Win Rate
Lane Change Acceleration Deceleration Average
PVG [9] 100.0% 90.5% 89.1% 93.2%
$\text{S}^{3}$aussian [28] 100.0% 97.9% 92.2% 96.7%
Deformable-GS [74] 95.8% 83.5% 72.9% 84.1%
Table 3: User study comparison of DriveDreamer4D win rates across various novel trajectory view synthesis.
4.2 Comparison with Different 4DGS Baselines
Quantitative Results. As demonstrated in Tab. 1, integrating DriveDreamer4D with different 4DGS algorithms consistently yields superior NTA-IoU and NTL-IoU scores across diverse, complex maneuvers (e.g., lane changes, acceleration, and deceleration), significantly outperforming the baseline methods. Specifically, with DriveDreamer4D, the average NTA-IoU scores across three baselines (PVG [9], $\text{S}^{3}$aussian [28], Deformable-GS [74]) are relatively enhanced by 22.6%, 43.5%, and 15.6%, underscoring DriveDreamer4D’s capability to improve the spatiotemporal coherence of foreground agents. Moreover DriveDreamer4D facilitates a relative improvement in average NTL-IoU for these baselines by 1.8%, 3.7%, and 1.6%, thereby markedly enhancing the spatiotemporal coherence of background lanes in 4D rendering of driving scenarios.
In addition to verifying the spatiotemporal consistency of rendered novel trajectory views, we leverage the FID metric to assess rendering quality under novel trajectories. Given that acceleration and deceleration scenarios yield rendered views with distributional similarities to ground truth, limiting FID’s discriminative capability across algorithms, our FID comparisons focus specifically on lane change scenarios. Experiment results, as presented in Tab. 2, indicate that our method substantially outperforms the baseline methods (PVG [9], $\text{S}^{3}$aussian [28], Deformable-GS [74]), with FID relative improvements of 32.1%, 46.4%, and 16.3%. These results highlight DriveDreamer4D’s capability to enhance generation quality for novel trajectory viewpoints.
Finally, we conduct a user study to evaluate the rendering quality of different methods on novel trajectories, with a specific focus on foreground agents. For each method, we generate three novel trajectory views—lane change, acceleration, and deceleration—across eight scenes from the Waymo dataset [57]. Participants are then asked to select the renderings they found most visually favorable in each comparison. The DriveDreamer4D win rates from this study, shown in Tab. 3, reveal a significant user preference for our method’s renderings.
$\lambda_{\text{novel}}$ NTA-IoU $\uparrow$ FID $\downarrow$
0 0.349 105.29
0.5 0.405 82.84
1 0.420 79.54
1.5 0.417 82.10
Table 4: Ablation study on the training loss weight $\lambda_{\text{novel}}$ for novel trajectory data.
$\lambda_{\text{reg}}$ NTA-IoU $\uparrow$ FID $\downarrow$
0 0.420 79.54
1e-2 0.411 119.39
1e-3 0.428 71.52
1e-4 0.422 75.31
Table 5: Ablation study on the regularization loss weight $\lambda_{\text{reg}}$.
Depth Loss Temporal-aligned Cousin Pair Regularization Loss NTA-IoU $\uparrow$ FID $\downarrow$
$\checkmark$ $\times$ $\times$ 0.401 82.63
$\times$ $\times$ $\times$ 0.420 79.54
$\times$ $\checkmark$ $\times$ 0.423 76.20
$\times$ $\checkmark$ $\checkmark$ 0.428 71.52
Table 6: Ablation study on the depth loss and CDTS.
Qualitative Results. In addition to quantitative comparisons, we provide a qualitative analysis of novel trajectory view renderings. As shown in Fig. 3, we present the novel trajectory view synthesis during lane change. Images rendered by the baseline algorithms exhibit issues where foreground vehicles incorrectly change lanes in sync with the camera’s motion, and some vehicles are incompletely rendered. Additionally, the background is filled with speckles and ghosting. Especially shown in the rightmost column of Fig. 3, baseline algorithms often produce blurred, ghosted foreground vehicles and background speckles in the sky, alongside blurred lane markings. Our method, however, significantly improves rendering quality, as highlighted by the orange boxes. Vehicle contours are sharper, and background artifacts such as speckles and ghosting are substantially reduced.
4.3 Ablation Studies
We conduct an ablation study based on PVG [9], analyzing the effects of hyperparameters settings for $\lambda_{\text{novel}}$ and $\lambda_{\text{reg}}$, as well as the impact of depth loss, temporal-aligned cousin pairs, and regularization loss. Following the experiments shown in Tab. 5 and 5, we set $\lambda_{\text{novel}}$ to 1 and $\lambda_{\text{reg}}$ to $1\times 10^{-3}$, which yield the best performance. As shown in Tab. 6, the experiment confirms that the depth loss should be excluded when optimizing novel trajectory views, since LiDAR depth maps are incomplete due to occlusions. Furthermore, employing temporal-aligned cousin pairs achieves an FID of 76.20 and an NTA-IoU of 0.423, with the perception regularization loss further improving the FID to 71.52 and NTA-IoU to 0.428. These results highlight the effectiveness of CDTS, in achieving $\sim$10improvement in FID, along with a $\sim$2increase in NTA-IoU.
5 Discussion and Conclusion
In this paper, we presented DriveDreamer4D, a novel framework designed to advance 4D driving scene representations by harnessing priors from world models. Addressing key limitations of current sensor simulation methods—namely, their dependence on forward-driving training data distributions and inability to model complex maneuvers—DriveDreamer4D leverages a world model to generate novel trajectory videos that complement real-world driving data. By explicitly employing structured conditions, our framework maintains spatial-temporal consistency across traffic elements, ensuring that generated data adheres closely to the dynamics of real-world traffic scenarios. Our experiments demonstrate that DriveDreamer4D achieves superior quality in generating diverse simulation viewpoints, with significant improvements in both the rendering fidelity and spatiotemporal coherence of scene components. Notably, these results highlight DriveDreamer4D’s potential as a foundation for closed-loop simulations that require high-fidelity reconstructions of dynamic driving scenes.